Q: ligjt travel at the speed of 300,000,000 m/s.
i) Express this speed in standard form
ii) Given that the mean distance from sun to Jupiter is 778.5 million kilometers, find the time taken in minutes and seconds, for light to travel from Sun to Jupiter.

Answer by Theo(13342)

the speed of light is 300,000,000 meters per second.

in scientific notation (also called standard form in england), this is equal to 3.0 * 10^8.

10^8 is equal to 100,000,000, therefore 3.0 * 10^8 is equal to 300,000,000.

the mean distance from the sun to jupiter is 778.5 million kilometers.

that would be 778.5 * 10^6 kilometers, in scientific notation.

rate * time = distance.

the rate of light is 3.0 * 10^8 kilometers per second.

the distance from the sun to jupiter is 778.5 * 10^6 kilometers.

the formula becomes 3.0 * 10^8 meters per second * time = 778.5 * 10^6 kilometers.

since the rate is in seconds, then time will be in seconds as well.

you need to translate the distance and speed into like measures.

you can either make them all kilometers or you can make them all meters.

i'll go with meters.

778.5 * 10^6 kilometers * 1000 = 778.5 * 10^6 * 10^3 = 778.5 * 10^9 meters.

the formula of rate * time = distance becomes:

3.0 * 10^8 * time = 778.5 * 10^9.

the speed is in meters per second and the distance is in meters and the time will be in seconds.

solve for time to get time = (778.5 * 10^9) / (3 * 10^8).

this can be shown as (778.5 / 3) * (10^9 / 10^8).

778.5 / 3 = 259.5
10^9 / 10^8 = 10^(9-8) = 10^1 = 10

formula for time becomes time = 259.5 * 10 = 2595 seconds.

there are 60 seconds per minute, therefore 2595 seconds / 60 = 43.25 minutes.

.25 minutes * 60 = 15 seconds.

therefore, the time it takes for light to travel from the sun to jupiter is 43 minutes and 15 seconds, assuming the distance is as given.

you have to make sure your measures are consistent with each other.

if you failed to see that the speed of light was per meter while the distance was kilometers, then the answer would not have been accurate.

i initially fell into that trap before i recognized my error.


it's all too easy to do.

i confirmed through a web search that the time it takes light to go from the sun to jupiter is right around what was calculated here, so the answer appears to be good.